# Stratified distance space improves the efficiency of sequential samplers for approximate Bayesian computation

### v1

@@ Abstract
Approximate Bayesian computation (ABC) methods are standard tools for inferring parameters of complex models when the likelihood function is analytically intractable. A popular approach to improving the poor acceptance rate of the basic rejection sampling ABC algorithm is to use sequential Monte Carlo (ABC SMC) to produce a sequence of proposal distributions adapting towards the posterior, instead of generating values from the prior distribution of the model parameters. Proposal distribution for the subsequent iteration is typically obtained from a weighted set of samples, often called particles, of the current iteration of this sequence. Current methods for constructing these proposal distributions treat all the particles equivalently, regardless of the corresponding value generated by the sampler, which may lead to inefficiency when propagating the information across iterations of the algorithm. To improve sampler efficiency, we introduce a modified approach called stratified distance ABC SMC. Our algorithm stratifies particles based on their distance between the corresponding synthetic and observed data, and then constructs distinct proposal distributions for all the strata. Taking into account the distribution of distances across the particle space leads to substantially improved acceptance rate of the rejection sampling. We further show that efficiency can be gained by introducing a novel stopping rule for the sequential process based on the stratified posterior samples and demonstrate these advances by several examples.
Keywords— Approximate Bayesian Computation, Importance Sampling, Likelihood-free inference, Sequential Monte Carlo
@@ 1 Introduction
Within the first extensions of ABC methods were sequential Monte Carlo ABC methods (Sisson et al., 2007; Toni et al., 2009; Moral et al., 2012). They improve on the efficiency of rejection sampling based ABC methods by introducing a sequence of importance sampling distributions to improve acceptance rate of the proposed simulated samples. ABC SMC methods are actively developed and many proposals have been made to improve the accuracy and efficiency (Silk et al., 2013; Filippi et al., 2013; Bonassi and West, 2015; Simola et al., 2020). Diverging from the sampling based ABC methods, synthetic likelihood (SL) methods (Wood, 2010; Price et al., 2018) directly find an estimate for the unknown likelihood function using a set of simulated samples. The development of both approaches, ABC and SL have been accelerated by the introduction of powerful machine learning techniques. Gutmann and Corander (2016) introduced an efficient Bayesian optimisation-based approach Bayesian optimisation for likelihood-free inference (BOLFI) to find a surrogate for the discrepancy metric as a function of the unknown parameters. Lueckmann et al. (2018); Papamakarios et al. (2019) introduced neural network-based flexible models that can be used as a surrogates for the likelihood function or directly for the posterior distribution.
@@ 2 Overview of ABC SMC methods
This basic ABC method is referred to as the rejection ABC. Arguably the main problem with it is the potentially poor acceptance rate associated with reasonable thresholds $\epsilon$ which is generally caused by the use of prior as a proposal distribution. To make ABC algorithm more efficient, importance sampling based sequential Monte Carlo ABC (ABC SMC) techniques were introduced to produce a sequentially improving set of proposal distributions. To improve the ABC posterior, the algorithm uses a decreasing sequence of acceptance thresholds $\epsilon_{1},\ldots,\epsilon_{T}$ and random walk proposal kernels $K_{t}(\theta^{(t+1)}\mid\theta^{(t)})$. Kernels are used to propagate a set of weighted samples $\left\{\left(\theta_{i}^{(t)},w_{i}^{(t)},y_{i}^{(t)}\right)\right\}_{i=1}^{N}$ that represent the current posterior distribution at iteration $t$ to $t+1$. The ABC posterior sample of the current iteration $t$ is often used to set the parameters of the kernel. Although other choices are possible, most often the kernels are zero-centered Gaussian distributions where the covariance matrices $\Sigma_{t}$ are empirically calculated from a set of weighted ABC posterior samples
@@
$q_{t}\left(\theta^{{(t)}},\theta^{(t+1)}\right)=\frac{p\left(\theta^{(t)}\mid y\in\Omega_{\left[0,\epsilon_{t}\right)}\right)K_{t}(\theta^{(t+1)}|\theta^{(t)})\int f(y\mid\theta^{(t+1)})\mathbb{I}\left(y\in\Omega_{\left[0,\epsilon_{t+1}\right)}\right)\mathsf{d}y}{\int\int p\left(\theta^{(t)}\mid y\in\Omega_{\left[0,\epsilon_{t}\right)}\right)K_{t}(\theta^{(t+1)}|\theta^{(t)})\int f(y\mid\theta^{(t+1)})\mathbb{I}\left(y\in\Omega_{\left[0,\epsilon_{t+1}\right)}\right)\mathsf{d}y\mathsf{d}\theta^{(t+1)}\mathsf{d}\theta^{(t)}}.$ (4)
@@
$q_{t}^{\prime}\left(\theta^{{(t)}},\theta^{(t+1)}\right)=p\left(\theta^{(t)}\mid y\in\Omega_{\left[0,\epsilon_{t}\right)}\right)p\left(\theta^{(t+1)}\mid y\in\Omega_{\left[0,\epsilon_{t+1}\right)}\right),$ (5)
@@
$\Sigma^{(t)}$ $=\int\int(\theta^{(t)}-\theta^{(t+1)})(\theta^{(t)}-\theta^{(t+1)})^{T}p\left(\theta^{(t)}\mid y\in\Omega_{\left[0,\epsilon_{t}\right)}\right)p\left(\theta^{(t+1)}\mid y\in\Omega_{\left[0,\epsilon_{t+1}\right)}\right)\mathsf{d}\theta^{(t)}\mathsf{d}\theta^{(t+1)},$ (7)
@@
$\left\{\left(\theta^{(t+1)}_{I(i)},\widetilde{w}^{(t+1)}_{I(i)},y^{(t+1)}_{I(i)}\right)\right\}_{i=1}^{N^{\prime}}=\left\{\left(\theta_{I(i)}^{(t)},\frac{w_{I(i)}^{(t)}}{\sum_{j=1}^{N^{\prime}}w_{I(j)}^{(t)}},y_{I(i)}^{(t)}\right):y_{I(i)}\in\Omega_{\left[0,\epsilon_{t+1}\right)}\right\}_{i=1}^{N^{\prime}}.$ (8)
@@
where the importance sampling distribution $\sum_{j=1}^{N}w_{j}^{(t)}K_{t}\left(\theta^{(t+1)}|\theta^{(t)}_{j}\right)$ is empirically estimated from the weighted sample as a Gaussian mixture.
@@
$\Sigma_{i}^{(t)}$ $=\int(\theta^{(t)}_{i}-\theta^{(t+1)})p\left(\theta^{(t+1)}\mid y\in\Omega_{\left[0,\epsilon_{t+1}\right)}\right)\mathsf{d}\theta^{(t+1)},$ (10)
@@ 3 Stratified sampling ABC SMC
In most cases ABC SMC is superior to the basic rejection ABC algorithm, especially when the prior distribution of the parameters is far off from the posterior distribution. There are multiple approaches for improving the standard ABC SMC, by choosing different proposal and target distributions, determining the threshold sequence adaptively or setting rules for stopping the sequential process. Especially in the early iterations of the algorithm, the kernels calculated empirically from the data can have large scale parameters which in turn can result in slower convergence to the approximate posterior distribution that is within our tolerance of error. We can investigate ABC SMC more closely by looking at how the particles are accepted and how the weights for the particles are being formed. At $t$th iteration, ABC posterior sample approximation is $\left\{\left(\theta_{i}^{(t)},w_{i}^{(t)},y_{i}^{(t)}\right)\right\}_{i=1}^{N}$. By using ABC SMC algorithm that was described in Section 2, the posterior approximation at $(t+1)$th iteration will be $\left\{\left(\theta_{i}^{(t+1)},w_{i}^{(t+1)},y_{i}^{(t+1)}\right)\right\}_{i=1}^{N_{\text{tot}}}$, where $\theta^{(t+1)}_{i}\stackrel{{\scriptstyle\text{i.i.d}}}{{\thicksim}}\sum_{j=1}^{N}w_{j}^{(t)}K_{t}(\theta^{(t+1)}\mid\theta_{j}^{(t)})$, $N_{\text{tot}}\geq N$ is the total number of simulated samples, and
@@
$w_{i}^{(t+1)}\propto\left\{\begin{array}[]{cc}\frac{p\left(\theta^{(t+1)}_{i}\right)}{\sum_{j=1}^{N}w_{j}^{(t)}K_{t}\left(\theta^{(t+1)}_{i}\mid\theta^{(t)}_{j}\right)},&\text{ if }y^{(t+1)}_{i}\in\Omega_{\left[0,\epsilon_{t+1}\right)},\left[\text{ w.p. }\mathsf{P}\left(y\in\Omega_{\left[0,\epsilon_{t+1}\right)}\mid\theta^{(t+1)}_{i}\right)\right]\\ 0,&\text{ if }y^{(t+1)}_{i}\notin\Omega_{\left[0,\epsilon_{t+1}\right)},\left[\text{ w.p. }\mathsf{P}\left(y\notin\Omega_{\left[0,\epsilon_{t+1}\right)}\mid\theta^{(t+1)}_{i}\right)\right]\end{array}\right.$ (13)
@@
It would be computationally desirable to generate as much samples as possible with non-zero weights, i.e. so that $y^{(t+1)}_{i}\in\Omega_{\left[0,\epsilon_{t+1}\right)}$. If $\mathsf{P}\left(y\in\Omega_{\left[0,\epsilon_{t+1}\right)}\mid\theta\right)$ would be available, it would be possible to query more samples from the parts of the parameter space that generate non-zero weighted samples with higher probability. Unfortunately this is in practice impossible in any realistic application of ABC. The average acceptance probability can be expressed as
$\int\int p(y\in\Omega_{\left[0,\epsilon_{t+1}\right)}\mid\theta^{(t+1)})K_{t}(\theta^{(t+1)}|\theta^{(t)})p\left(\theta^{(t)}\mid y\in\Omega_{\left[0,\epsilon_{t}\right)}\right)\mathsf{d}\theta^{(t)}\mathsf{d}\theta^{(t+1)},$ (14)
@@
which we aim to improve especially at the early iteration rounds $t$ by introducing a novel importance sampling approach that modifies the locally optimal ABC SMC defined by the kernel covariance (10) using a partition of the ABC posterior approximation into components defined by a set of non-overlapping acceptance regions. The observation space bounded to the acceptance region at first threshold of ABC SMC can be partitioned based on the threshold sequence
$\Omega_{\left[0,\epsilon_{1}\right)}=\bigcup_{t=1}^{T}\Omega_{\left[\epsilon_{t+1},\epsilon_{t}\right)},$ (15)
@@
$p\left(\theta^{(t)}\mid y\in\Omega_{\left[0,\epsilon_{t}\right)}\right)$ $=\sum_{k=t}^{T}p\left(\theta^{(t)}\mid y\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}\right)\mathsf{P}\left(y\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}\mid y\in\Omega_{\left[0,\epsilon_{t}\right)}\right).$ (16)
@@
Figure 1: Example 1. Partitioning of the ABC posterior distribution based on the acceptance regions.
@@
The partitioned posterior motivates the proposed ABC SMC approach. We make a modification to the transition kernel by removing the dependence on iteration $t$, and instead define it for each posterior component conditioned on the events of observations belonging to the acceptance regions $\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}$. Given the set of predefined acceptance regions and the ABC posterior sample at $i$th iteration, we first define a locally optimal transition kernel similar to (10) by taking into account which partitioned acceptance region band a particle resides in and then balance the particle selection probability by using approximations to the weights in (16) that we get as a by-product of previous iterations of ABC SMC.
@@
Let $(\theta_{i}^{(t)},y_{i}^{(t)},w_{i}^{(t)})\in\left\{(\theta^{(t)},y^{(t)},w^{(t)})\mid y^{(t)}\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}\right\}$ be a tuple that defines a particle from ABC posterior at iteration $t\leq k<T$, that is within a partitioned acceptance region band $\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}$. We seek a transition kernel covariance that is optimal for mutating the particle into acceptance region $\Omega_{\left[0,\epsilon_{k+1}\right)}$. This can be achieved using the same approach as deriving locally optimal transition kernel covariance (10)
@@
$\Sigma_{i}^{(k)}$ $=\int\left(\theta-\theta_{i}^{(t)}\right)\left(\theta-\theta_{i}^{(t)}\right)^{T}p\left(\theta\mid y\in\Omega_{\left[0,\epsilon_{k+1}\right)}\right)\mathsf{d}\theta.$
@@
In Figure 2 we illustrate the cascading nature of the ABC posterior approximations at a single iteration $t$ of the methods and how the transition kernels are locally optimally defined for each of the particles separately based on posterior component they belong to.
@@
We express the average acceptance rate (14) as
$\int\int\sum_{l=t+1}^{T}\sum_{k=t}^{T}p(y\in\Omega_{\left[\epsilon_{l+1},\epsilon_{l}\right)}\mid\theta^{(t+1)})K_{k}(\theta^{(t+1)}|\theta^{(t)})p\left(\theta^{(t)}\mid y\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}\right)\mathsf{d}\theta^{(t)}\mathsf{d}\theta^{(t+1)}$
$=\sum_{l=t+1}^{T}\sum_{k=t}^{T}\int\int p(y\in\Omega_{\left[\epsilon_{l+1},\epsilon_{l}\right)}\mid\theta^{(t+1)})K_{k}(\theta^{(t+1)}|\theta^{(t)})p\left(\theta^{(t)}\mid y\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}\right)\mathsf{d}\theta^{(t)}\mathsf{d}\theta^{(t+1)}$
@@
where $C_{l,k}^{(t)}$ is the average rate of transitioning a particle from posterior $p(\theta^{(t)}\mid y\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)})$ with kernel $K_{k}(\cdot\mid\cdot)$ and generating a synthetic observation $y\in\Omega_{\left[\epsilon_{l+1},\epsilon_{l}\right)}$. Given that we are approximating $C_{l,k}^{(t)}\approx C_{l,k},l=t+1,\ldots,T,k=t,\ldots,T$ at each iteration of the ABC SMC, we can use these regional weights as a basis for a stratification strategy to compose an importance sampling distribution for more efficient simulation acceptance, by drawing samples from different partitioned posterior components proportionally to their average acceptance rates $C_{l,k}$. In practice we approximate $C_{l,k}$ by counting the frequencies $f_{l,k}^{(t)}$ how often samples drawn from

### v2

@@ Abstract
Approximate Bayesian computation (ABC) methods are standard tools for inferring parameters of complex models when the likelihood function is analytically intractable. A popular approach to improving the poor acceptance rate of the basic rejection sampling ABC algorithm is to use sequential Monte Carlo (ABC SMC) to produce a sequence of proposal distributions adapting towards the posterior, instead of generating values from the prior distribution of the model parameters. Proposal distribution for the subsequent iteration is typically obtained from a weighted set of samples, often called particles, of the current iteration of this sequence. Current methods for constructing these proposal distributions treat all the particles equivalently, regardless of the corresponding value generated by the sampler, which may lead to inefficiency when propagating the information across iterations of the algorithm. To improve sampler efficiency, a modified approach called stratified distance ABC SMC is introduced. The algorithm stratifies particles based on their distance between the corresponding synthetic and observed data, and then constructs distinct proposal distributions for all the strata. Taking into account the distribution of distances across the particle space leads to substantially improved acceptance rate of the rejection sampling. It is shown that further efficiency could be gained by using a newly proposed stopping rule for the sequential process based on the stratified posterior samples and these advances are demonstrated by several examples.
@@ 1 Introduction
Within the first extensions of ABC methods were sequential Monte Carlo ABC methods (Sisson et al., 2007; Toni et al., 2009; Moral et al., 2012). They improve on the efficiency of rejection sampling based ABC methods by introducing a sequence of importance sampling distributions to improve acceptance rate of the proposed simulated samples. ABC SMC methods are actively developed and many proposals have been made to improve the accuracy and efficiency by introducing improved importance sampling distribution generation strategies or adaptive threshold selection (Silk et al., 2013; Filippi et al., 2013; Bonassi and West, 2015; Simola et al., 2020). Other strategies to improve sequential sampling is to incorporate either multilevel Monte Carlo strategies (Jasra et al., 2019) or delayed acceptance sampling (Everitt and Rowińska, 2021) where computationally faster lower level approximate models can be used to reduce the number of generated data points from the expensive accurate models. Diverging from the sampling based ABC methods, synthetic likelihood (SL) methods (Wood, 2010; Price et al., 2018) directly find an estimate for the unknown likelihood function using a set of simulated samples. The development of both approaches, ABC and SL have been accelerated by the introduction of powerful machine learning techniques. Gutmann and Corander (2016) introduced an efficient Bayesian optimisation-based approach Bayesian optimisation for likelihood-free inference (BOLFI) to find a surrogate for the discrepancy metric as a function of the unknown parameters. Lueckmann et al. (2018) and Papamakarios et al. (2019) introduced neural network-based flexible models that can be used as a surrogates for the likelihood function or directly for the posterior distribution.
@@ 2 Overview of ABC SMC methods
This basic ABC method is referred to as the rejection ABC. Arguably the main problem with it is the potentially poor acceptance rate associated with reasonable thresholds $\epsilon$ which is generally caused by the use of prior as a proposal distribution. To make ABC algorithm more efficient, importance sampling based sequential Monte Carlo techniques (Chopin and Papaspiliopoulos, 2020) have been applied in ABC. The resulting sequential Monte Carlo ABC (ABC SMC) techniques were introduced to produce a sequentially improving set of proposal distributions. To improve the ABC posterior, the algorithm uses a decreasing sequence of acceptance thresholds $\epsilon_{1},\ldots,\epsilon_{T}$ and random walk proposal kernels $K_{t}(\theta^{(t+1)}\mid\theta^{(t)})$. Kernels are used to propagate a set of weighted samples $\left\{\left(\theta_{i}^{(t)},w_{i}^{(t)},y_{i}^{(t)}\right)\right\}_{i=1}^{N}$ that represent the current posterior distribution at iteration $t$ to $t+1$. The ABC posterior sample of the current iteration $t$ is often used to set the parameters of the kernel. Although other choices are possible, most often the kernels are zero-centered Gaussian distributions where the covariance matrices $\Sigma_{t}$ are empirically calculated from a set of weighted ABC posterior samples
@@
$q_{t}\left(\theta^{{(t)}},\theta^{(t+1)}\right)$ (4)
$=\frac{p\left(\theta^{(t)}\mid y\in\Omega_{\left[0,\epsilon_{t}\right)}\right)K_{t}(\theta^{(t+1)}|\theta^{(t)})\mathsf{P}\left(y^{\prime}\in\Omega_{\left[0,\epsilon_{t+1}\right)}\mid\theta^{(t+1)}\right)}{\int\int p\left(\theta^{(t)}\mid y\in\Omega_{\left[0,\epsilon_{t}\right)}\right)K_{t}(\theta^{(t+1)}|\theta^{(t)})\mathsf{P}\left(y^{\prime}\in\Omega_{\left[0,\epsilon_{t+1}\right)}\mid\theta^{(t+1)}\right)\mathsf{d}\theta^{(t+1)}\mathsf{d}\theta^{(t)}}.$
@@
$q_{t}^{\prime}\left(\theta^{{(t)}},\theta^{(t+1)}\right)=p\left(\theta^{(t)}\mid y\in\Omega_{[0,\epsilon_{t})}\right)p\left(\theta^{(t+1)}\mid y^{\prime}\in\Omega_{[0,\epsilon_{t+1})}\right)$ (5)
@@
$\Sigma^{(t)}=\int\!\!\int$ $(\theta^{(t)}\!-\!\theta^{(t+1)})(\theta^{(t)}\!-\!\theta^{(t+1)})^{T}\times$
×$\times p\left(\theta^{(t)}\mid y\in\Omega_{\left[0,\epsilon_{t}\right)}\right)p\left(\theta^{(t+1)}\mid y^{\prime}\in\Omega_{[0,\epsilon_{t+1})}\right)\mathsf{d}\theta^{(t)}\mathsf{d}\theta^{(t+1)},$ (7)
@@
$\left\{\left(\theta^{(t+1)}_{I(i)},\widetilde{w}^{(t+1)}_{I(i)},y^{(t+1)}_{I(i)}\right)\right\}_{i=1}^{N^{\prime}}=\left\{\left(\theta_{I(i)}^{(t)},\frac{w_{I(i)}^{(t)}}{\sum_{j=1}^{N^{\prime}}w_{I(j)}^{(t)}},y_{I(i)}^{(t)}\right):y_{I(i)}\in\Omega_{\left[0,\epsilon_{t+1}\right)}\right\}_{i=1}^{N^{\prime}},$ (8)
where the subset of indices $\{I(i)\}_{i=1}^{N^{\prime}}\subset\{1,2,\ldots,N\}$ indicate samples for which $y_{I(i)}\in\Omega_{\left[0,\epsilon_{t+1}\right)}$.
@@
where $p(\cdot)$ is the prior of the parameters and the importance sampling distribution $\sum_{j=1}^{N}w_{j}^{(t)}K_{t}\left(\theta^{(t+1)}|\theta^{(t)}_{j}\right)$ is empirically estimated from the weighted sample as a Gaussian mixture.
@@
$\Sigma_{i}^{(t)}$ $=\int(\theta^{(t)}_{i}-\theta^{(t+1)})(\theta^{(t)}_{i}-\theta^{(t+1)})^{T}p\left(\theta^{(t+1)}\mid y\in\Omega_{\left[0,\epsilon_{t+1}\right)}\right)\mathsf{d}\theta^{(t+1)},$ (10)
@@ 3 Stratified sampling ABC SMC
In most cases ABC SMC is superior to the basic rejection ABC algorithm, especially when the prior distribution of the parameters is far off from the posterior distribution. There are multiple approaches for improving the standard ABC SMC, by choosing different proposal and target distributions, determining the threshold sequence adaptively or setting rules for stopping the sequential process. Especially in the early iterations of the algorithm, the kernels calculated empirically from the data can have large scale parameters which in turn can result in slower convergence to the approximate posterior distribution that is within our tolerance of error. We can investigate ABC SMC more closely by looking at how the particles are accepted and how the weights for the particles are being formed. At $t$th iteration, ABC posterior sample approximation is $\left\{\left(\theta_{i}^{(t)},w_{i}^{(t)},y_{i}^{(t)}\right)\right\}_{i=1}^{N}$. By using ABC SMC algorithm that was described in Section 2, the posterior approximation at $(t+1)$th iteration will be $\left\{\left(\theta_{i}^{(t+1)},w_{i}^{(t+1)},y_{i}^{(t+1)}\right)\right\}_{i=1}^{N_{\text{tot}}^{(t+1)}}$, where $\theta^{(t+1)}_{i}\mid\theta_{j}^{(t)}\stackrel{{\scriptstyle\text{i.i.d}}}{{\thicksim}}\sum_{j=1}^{N}w_{j}^{(t)}K_{t}(\theta^{(t+1)}\mid\theta_{j}^{(t)})$, $N_{\text{tot}}^{(t+1)}\geq N$ is the total number of simulated samples at $(t+1)$th iteration, and
@@
$w_{i}^{(t+1)}\propto\left\{\begin{array}[]{cc}\frac{p\left(\theta^{(t+1)}_{i}\right)}{\sum_{j=1}^{N}w_{j}^{(t)}K_{t}\left(\theta^{(t+1)}_{i}\mid\theta^{(t)}_{j}\right)},&\text{ if }y^{(t+1)}_{i}\in\Omega_{\left[0,\epsilon_{t+1}\right)},\left[\text{ w.p. }\mathsf{P}\left(y\in\Omega_{\left[0,\epsilon_{t+1}\right)}\mid\theta^{(t+1)}_{i}\right)\right]\\ 0,&\text{ if }y^{(t+1)}_{i}\notin\Omega_{\left[0,\epsilon_{t+1}\right)},\left[\text{ w.p. }\mathsf{P}\left(y\notin\Omega_{\left[0,\epsilon_{t+1}\right)}\mid\theta^{(t+1)}_{i}\right)\right].\end{array}\right.$ (13)
@@
Samples are generated until $N$ samples have non-zero weights, so it would be computationally desirable to generate as many samples as possible so that $y^{(t+1)}_{i}\in\Omega_{\left[0,\epsilon_{t+1}\right)}$, and achieve a large acceptance rate $N/N^{(t+1)}_{\text{tot}}$. If $\mathsf{P}\left(y\in\Omega_{\left[0,\epsilon_{t+1}\right)}\mid\theta\right)$ would be available, it would be possible to query more samples from the parts of the parameter space that generate non-zero weighted samples with higher probability. Unfortunately this is in practice impossible in any realistic application of ABC. The average acceptance probability at $(t+1)$th iteration can be expressed as
$\int\int\mathsf{P}\left(y^{\prime}\in\Omega_{\left[0,\epsilon_{t+1}\right)}\mid\theta^{(t+1)}\right)K_{t}(\theta^{(t+1)}|\theta^{(t)})p\left(\theta^{(t)}\mid y\in\Omega_{\left[0,\epsilon_{t}\right)}\right)\mathsf{d}\theta^{(t)}\mathsf{d}\theta^{(t+1)},$ (14)
@@
which we aim to improve especially at the early iteration rounds $t$ by introducing a novel importance sampling approach that modifies the locally optimal ABC SMC defined by the kernel covariance (10) using a partition of the ABC posterior approximation into components defined by a set of non-overlapping acceptance regions. The observation space bounded to the acceptance region at threshold $\epsilon_{t}$ of ABC SMC can be partitioned based on the threshold sequence
$\Omega_{\left[0,\epsilon_{t}\right)}=\bigcup_{k=t}^{T}\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)},$ (15)
@@
$p\left(\theta^{(t)}\mid y\in\Omega_{\left[0,\epsilon_{t}\right)}\right)$ $=\sum_{k=t}^{T}p(\theta^{(t)}\mid y^{\prime}\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)})\mathsf{P}\left(y^{\prime}\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}\mid y\in\Omega_{\left[0,\epsilon_{t}\right)}\right).$ (16)
@@
Figure 1: Example 1. Partitioning of the ABC posterior distribution based on the acceptance regions. Here the posterior is actually the prior as the acceptance region $\Omega_{0,\infty}$ is the whole observation space. In the first round of ABC SMC we’d wish to transfer particles $p(\theta\mid y\in\Omega_{0,\infty})$ into $p(\theta\mid y\in\Omega_{0,4})$ that is a combination of the four other posterior components.
@@
The partitioned posterior motivates the proposed ABC SMC approach. We propose a modification to the locally optimal Gaussian transition kernel by removing the dependence on iteration $t$, and instead define it for each posterior component conditioned on the events of observations belonging to the acceptance regions $\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}$. Given the set of predefined acceptance regions and the ABC posterior sample at $i$th iteration, we first define a locally optimal transition kernel similar to (10) by taking into account which partitioned acceptance region band a particle resides in and then balance the particle selection probability by using approximations to the weights in (16) that we get as a by-product of previous iterations of ABC SMC.
@@
Let $(\theta_{i}^{(t)},y_{i}^{(t)},w_{i}^{(t)})\in\left\{(\theta^{(t)},y^{(t)},w^{(t)})\mid y^{(t)}\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}\right\}$ be a tuple that defines a particle from ABC posterior at iteration $t\leq k<T$, that is within a partitioned acceptance region band $\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}$. We seek a transition kernel
$K_{k}(\theta^{(t+1)}\mid\theta^{(t)}_{i})=\mathsf{Normal}\left(\theta^{(t+1)}\mid\theta^{(t)},\Sigma_{i}^{(k)}\right)$ (18)
where covariance $\Sigma_{i}^{(k)}$ is optimal for mutating the particle $\theta^{(t)}_{i}$ associated with $\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}$ so that it would produce an observation from acceptance region $\Omega_{\left[0,\epsilon_{k+1}\right)}$. The idea here is that the particles from each iteration $t$ that are associated with acceptance region $\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}$ would the transferred to a posterior that is defined by $\Omega_{\left[0,\epsilon_{k+1}\right)}$ that is one step closer to the final acceptance region $\Omega_{\left[0,\epsilon_{T}\right)}$ as defined by the fixed threshold sequence. Traditionally, all particles would be transferred to approximate a posterior defined by $\Omega_{\left[0,\epsilon_{t+1}\right)}$, regardless of the associated values $y_{i}^{(t)}$. The proposed modification can be achieved using the same approach as deriving locally optimal transition kernel covariance (10). Now proposed ABC SMC joint proposal distribution for the tranformation at $t$th iteration is of the form
$q_{k}\left(\theta^{{(t)}},\theta^{(t+1)}\right)$ (19)
@@
$=\frac{p\left(\theta^{(t)}\mid y\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}\right)K_{k}(\theta^{(t+1)}|\theta^{(t)})\mathsf{P}\left(y^{\prime}\in\Omega_{\left[0,\epsilon_{k+1}\right)}\mid\theta^{(t+1)}\right)}{\int\int p\left(\theta^{(t)}\mid y\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}\right)K_{k}(\theta^{(t+1)}|\theta^{(t)})\mathsf{P}\left(y^{\prime}\in\Omega_{\left[0,\epsilon_{k+1}\right)}\mid\theta^{(t+1)}\right)\mathsf{d}\theta^{(t+1)}\mathsf{d}\theta^{(t)}}.$
As before, we derive an optimal Gaussian kernel (3) by minimizing the Kullback-Leibler (KL) discrepancy between (4) and
$q_{k}^{\prime}\left(\theta^{(t)},\theta^{(t+1)}\right)=p\left(\theta^{(t)}\mid y\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}\right)p\left(\theta^{(t+1)}\mid y^{\prime}\in\Omega_{\left[0,\epsilon_{k+1}\right)}\right),$ (20)
which leads to
$\Sigma^{(k)}=\int\!\!\int$ $(\theta^{(t)}\!-\!\theta^{(t+1)})(\theta^{(t)}\!-\!\theta^{(t+1)})^{T}\times$
×$\times p\left(\theta^{(t)}\mid y\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}\right)p\left(\theta^{(t+1)}\mid y^{\prime}\in\Omega_{\left[0,\epsilon_{k+1}\right)}\right)\mathsf{d}\theta^{(t)}\mathsf{d}\theta^{(t+1)},$ (21)
for which a locally optimal version at $\theta^{(t)}_{i}$ is
$\Sigma_{i}^{(k)}$ $=\int\left(\theta_{i}^{(t)}-\theta^{(t+1)}\right)\left(\theta_{i}^{(t)}-\theta^{(t+1)}\right)^{T}p\left(\theta^{(t+1)}\mid y\in\Omega_{\left[0,\epsilon_{k+1}\right)}\right)\mathsf{d}\theta,$ (22)
@@
where $\theta_{i}^{(t)}$ is such that the associated observation $y_{i}^{(t)}\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}$. In Figure 2 we illustrate the cascading nature of the ABC posterior approximations at a single iteration $t$ of the methods and how the transition kernels are locally optimally defined for each of the particles separately based on posterior component they belong to.
@@
We express the average acceptance rate at the transition from round $t\rightarrow t+1$ (14) as
$\int\int\sum_{l=t+1}^{T}\sum_{k=t}^{T}\mathsf{P}\left(y^{\prime}\in\Omega_{\left[\epsilon_{l+1},\epsilon_{l}\right)}\mid\theta^{(t+1)}\right)K_{k}(\theta^{(t+1)}|\theta^{(t)})p\left(\theta^{(t)}\mid y\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}\right)\mathsf{d}\theta^{(t)}\mathsf{d}\theta^{(t+1)}$
$=\sum_{l=t+1}^{T}\sum_{k=t}^{T}\int\int\mathsf{P}\left(y^{\prime}\in\Omega_{\left[\epsilon_{l+1},\epsilon_{l}\right)}\mid\theta^{(t+1)}\right)K_{k}(\theta^{(t+1)}|\theta^{(t)})p\left(\theta^{(t)}\mid y\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}\right)\mathsf{d}\theta^{(t)}\mathsf{d}\theta^{(t+1)}$
@@
where $C_{l,k}^{(t)}$ is the average rate of transitioning a particle from posterior $p(\theta^{(t)}\mid y\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)})$ with kernel $K_{k}(\cdot\mid\cdot)$ and generating a synthetic observation $y\in\Omega_{\left[\epsilon_{l+1},\epsilon_{l}\right)}$. Given that we are approximating $C_{l,k}^{(t)}\approx C_{l,k},l=t+1,\ldots,T,k=t,\ldots,T$ at each iteration of the ABC SMC, we can use these regional weights as a basis for a stratification strategy to compose an importance sampling distribution for more efficient simulation acceptance, by using the information we have obtained about the different different partitioned posterior components and their average acceptance rates $C_{l,k}$. In practice we approximate $C_{l,k}$ by counting the frequencies $f_{l,k}^{(t)}$ how often samples drawn from